OpenStudy (anonymous):

Solve the linear equations: 159x + 3.3y = 1000 2.3x + 0.6y = 18

OpenStudy (anonymous):

Start by multiplying both equations by 10 to get rid of that decimal.

OpenStudy (anonymous):

SOlution\[x = 6.15647\]\[y = 6.4\]

OpenStudy (anonymous):

@mattpyne Please avoid giving straight solutions and instead helping explain the problems.

OpenStudy (anonymous):

what do you mean by bet rid of the decimal?

OpenStudy (anonymous):

Well, it is hard to solve the equation with decimals in it, at least without a calculator. So instead let's multiply both equations by 10 to move the decimal one place to the right.

OpenStudy (anonymous):

What would you get?

OpenStudy (anonymous):

oh okay 1590x + 33y = 1000 23x + 6y = 18

OpenStudy (anonymous):

oh I guess 1000 would be 10,000 and 18 would be 180

OpenStudy (anonymous):

Yes. Now we have two methods to solve this equation. We can use elimination or substitution. Elimination means making one of the variables the same as another. For example) 2x+6y=4 3x+3y=2 or 6x+6y Subtract the two equations. Now we have: 2x+6y=4 -(6x+6y=0) -------- -4x=4 x=-1, plug in x to solve for y.

OpenStudy (anonymous):

I don't quite understand how you got these equations...

OpenStudy (anonymous):

Sorry, made a mistake in the above equation, should be 6x+6y=4.

OpenStudy (anonymous):

Okay, so these are called system of equation problems. This means the x and y values in both equations are the same. We have two methods, and I'm going to make it simple by suggesting elimination for this one. As you said, after getting rid of the decimals we have: 1590x + 33y = 10000 23x + 6y = 180

OpenStudy (anonymous):

Let us find the LCM or least common multiple of 33y and 6y. For example, 4 and 6 have an LCM of 12 because 4*3=12 and 6*2=12. 4 and 6 multiplied by some number both equal 12.

OpenStudy (anonymous):

Try to find something that is divisible by both 33 and 6.

OpenStudy (anonymous):

3.... 33/3=11 6/3=2

OpenStudy (anonymous):

Think of what you did, but the opposite. You want a number x that when x/33 you get any number and when x/6 you get another number.

OpenStudy (anonymous):

I do not get it....any number I multiply gets a tiny number....